### v4

# Constitutional Consensus

Abstract
Consider people with smartphones operating without external authorities or global resources other than the network itself. In this setting, high-end applications supporting sovereign democratic digital communities, community banks, and digital cooperatives require consensus executed by community members, which must be reconfigurable to support community dynamics.
The Constitutional Consensus protocol aims to address this need by introducing constitutional self-governance to consensus: participants dynamically amend the participant set, supermajority threshold, and timeout parameter through the consensus protocol itself. We achieve this by enhancing a DAG-based protocol (like Cordial Miners) with participant-controlled reconfiguration, while also supporting both high- and low-throughput operation (like Morpheus), remaining quiescent when idle. This three-way synthesis uniquely combines: (1) constitutional amendments for self-governance, (2) a cryptographic DAG structure for simplicity, parallelism, and throughput, and (3) both high- and low-throughput operation. The protocol achieves consensus in $3\delta$, maintains $O(n)$ amortized communication complexity during high throughput, and seamlessly transitions between modes. The basic protocol (without constitutional amendments) realizes these features in 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
keywords:
Consensus Protocols, Grassroots Platforms, Blocklace, Digital Social Contracts, Democratic Digital Communities
1 Introduction
Participant-Controlled Consensus. Consider people who wish to form self-governing digital communities without relying on servers or external authorities. These participants—whether members of community banks, digital cooperatives, or democratic assemblies—need consensus protocols they can execute on their smartphones while maintaining complete control over membership and governance rules. Unlike traditional reconfiguration where external authorities determine changes [40, 2], these communities require constitutional self-governance: participants must collectively control who joins, how decisions are made, and how the protocol operates, all through the consensus protocol itself. This paper presents Constitutional Consensus, a protocol that enables such self-governance by allowing participants to dynamically amend the participant set $P$, supermajority threshold $\sigma$, and timeout parameter $\Delta$ through democratic decisions implemented within the protocol.
Digital Social Contracts. These self-governing communities implement digital social contracts [9]—deterministic state machines that encode agreements among people who know each other. Like smart contracts, they provide computational guarantees about contract execution. However, while smart contracts [10] operate among anonymous accounts and run on consensus protocols executed by third-party servers, social contracts are among identified people and run on consensus protocols executed by the participants’ own devices. Existing consensus-based smart contracts, notably in Decentralized Finance (DeFi) platforms [10, 7, 25], have driven development of high-performance consensus protocols [43, 17, 18, 4, 16], but these operate among pseudonymous accounts on global networks rather than identified people in local communities. This fundamental difference enables democratic rather than plutocratic governance, as shown in Table 1.
Table 1: Comparing Smart Contracts and Consensus-Based Digital Social Contracts
Smart Contract [10] Digital Social Contract [9]
Among: Accounts People
Executed by: Third-party servers The people’s smartphones
In return for: Gas Naught
Applications: Decentralized Finance (DeFi) [31, 15],
Decentralized Autonomous Organizations (DAOs) [12, 11],
Non-Fungible Tokens (NFTs) [41]
Grassroots democratic digital communities [28, 33] and their federation [38, 39],
community banks and digital cooperatives [37]
Amended via: Self-modifying code Constitutional amendments
Governance: Token-weighted voting (plutocracy) One person, one vote (democracy)
Amendment Authority: External governance tokens or DAOs Participants themselves via supermajority
Parameter Control: Fixed in code or controlled by token holders Democratically adjustable by participants
We note that while smart contracts are necessarily consensus-based, digital social contracts are not. Some grassroots platforms, notably grassroots social networks [35] and grassroots cryptocurrencies [37, 22], do not require consensus on the total ordering of the acts of their participants: Their safety depends only on the agents obeying the social contract [38]. This can be ensured by mechanisms weaker then consensus, e.g. mutual attestations among agents operating within Trusted Execution Environments (TEEs) [30]. However, applications requiring collective agreement—such as community banks with multiple signatories or democratic governance with voting—do require consensus to ensure all participants agree on the order of transactions and decisions.
Technical Requirements. Communities implementing digital social contracts face unique technical challenges. Small communities may have transaction rates much lower than network latency, requiring a protocol that remains quiescent—sending no messages when idle. Yet the same protocol must efficiently handle high-throughput periods when the community is larger and/or more active. During low throughput, transactions arrive one-at-a-time with no conflicts to resolve, so ideally any participant should be able to finalize their transaction without designated leaders. During high throughput, the protocol should match the performance of modern consensus protocols.
Specifically, an effective protocol for self-governing communities requires:
  1. 1.
    Quiescent operation: No activity when idle, spontaneous leader emergence for isolated transactions
  2. 2.
    Dual-mode efficiency: Seamless transition between low and high throughput modes
  3. 3.
    Responsiveness: Network-speed operation unaffected by conservative timeout estimates
  4. 4.
    DAG-based structure: Leveraging blocklace [3] parallelism and integration
  5. 5.
    Constitutional amendments: Participant-controlled governance of protocol parameters
Technical Foundation. Constitutional Consensus achieves these requirements through a three-way synthesis. From Morpheus [23], it inherits dual-mode operation with spontaneous leaders during low throughput and formal leadership during high throughput. From Cordial Miners [18], it adopts the blocklace structure for parallelism and efficiency. To these, it adds constitutional amendment mechanisms enabling true self-governance. The protocol operates in waves of three rounds. During low throughput, any participant with a transaction can spontaneously lead a wave, achieving finality in $3\delta$. During high throughput, formal leaders are selected round-robin, maintaining $O(n)$ amortized communication complexity. The basic protocol (without amendments) requires only 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
Constitutional Self-Governance. The key innovation lies in how participants control their own governance. Unlike smart contracts that rely on external governance tokens or DAOs [12], constitutional amendments are decided by the participants themselves through supermajority vote. While Lamport’s original Paxos envisioned a self-governing parliament [21], subsequent reconfiguration protocols [42, 2, 40, 20] assume external authorities determine membership changes. Constitutional Consensus returns to Paxos’s original vision of true self-governance. Unlike Proof-of-Stake systems where validators must stake capital [8], participation requires only the democratic approval of existing members. This creates a fundamental architectural difference: while blockchain systems typically support only a single global instance whose security depends on token value, Constitutional Consensus enables unlimited independent communities to self-govern simultaneously.
Connection to Grassroots Architecture. This capability for multiple independent instances operating without central coordination or global resources makes Constitutional Consensus inherently grassroots [38]. A community in Bangladesh, a cooperative in Brazil, and a federation in Berlin can each run their own instance, potentially with overlapping members, without any coordination or competition for resources. This architecture—multiple local instances rather than a single global instance—enables the broader grassroots vision of egalitarian digital platforms that operate without central authorities [36, 34]. Within this grassroots architecture, Constitutional Consensus provides the consensus layer for applications requiring collective agreement, complementing consensus-free grassroots platforms like social networks [35] and cryptocurrencies [37, 22]. Together, these protocols enable communities to build sovereign democratic digital infrastructure from the ground up, controlled entirely by their members rather than external authorities or token holders. This distinguishes Constitutional Consensus from global platforms—whether centralized like Facebook or decentralized like Bitcoin [26] and IPFS [5]—that can support only a single instance.
Example: Grassroots Communities, Banks, and Federations. As a concrete scenario for grassroots platforms, consider a small town whose residents wish to form a digital community. They begin by deploying independent instances of consensus-free grassroots platforms on their smartphones—creating their own grassroots social network [35] and trading via self-minted grassroots cryptocurrencies [37, 22]. Crucially, they do this without coordination with any global network or central authority. When their needs evolve to require collective agreement—community banks with multiple signatories, constitutional governance, or federation with neighbouring towns—they deploy their own Constitutional Consensus instance. This instance operates independently yet can interoperate with other communities’ instances, embodying the grassroots principle of multiple local instances rather than a single global system. Through Grassroots Federation [39], these autonomous instances can scale democratically from local to potentially global governance. Small communities ($\approx 100$ members) govern themselves directly; larger ones use sortition-selected assemblies [14, 39]. Each community’s governance runs as a digital social contract on their own Constitutional Consensus instance, executed by members on their smartphones—no external servers or authorities required.
Model. As deterministic fault-tolerant consensus is not solvable in asynchronous systems [13], we assume an eventual synchrony model, where communication delays are at most $\delta$ after some Global Stabilization Time (GST) that is unknown to the agents. We consider Byzantine failures, as some malicious actors may infiltrate the network, but we assume a bounded ratio of Byzantine agents (to be specified later), since, by and large, the agents trust each other.
Paper structure. Section 2 provides the model and problem definitions. Section 3 presents a single-epoch Constitutional Consensus protocol, with only an initial constitution and no constitutional amendments. Section 4 proves the safety and liveness of single-epoch Constitutional Consensus. Section 5 adds support for constitutional amendments – for each of the three parameters $P,\sigma$, and $\Delta$, we present a democratic process for amending it. We then present the protocol extension needed to realize constitutional amendments and prove its safety and liveness. Section 6 proves that Constitutional Consensus is grassroots. Section 7 concludes.
2 Model and Problem
Model. We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$. While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it. A correct agent follows the protocol. An agent that does not follow the protocol is faulty or Byzantine.
We assume an eventual synchrony (aka partial synchrony) model in which every message sent between correct agents is eventually received. Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
Note that the abstraction of reliable messaging does not mean that agents must communicate via point-to-point TCP connections. Rather, it abstracts away a dissemination layer that can be implemented using acks/nacks, gossip, network coding, etc., and is orthogonal to the consensus protocol in this paper.
A protocol proceeds through a sequence of epochs, each commences with a new constitution (defined next), referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution. The constitution defines the set of participants in the epoch, as well as key protocol parameters:
Definition 2.1 (Constitution, $\sigma$-Supermajority).
A constitution is a triple $(P,\sigma,\Delta)$ where:
  1. 1.
    $P\subset\Pi$ is a set of agents, with $n:=|P|$,
  2. 2.
    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the agents $P$,
  3. 3.
    $\Delta>0$ is a presumed upper-bound on message delay after GST among $P$.
We may omit $\Delta$ when its value is immaterial (e.g., when discussing protocol safety) and refer to the constitution $(P,\sigma)$. A $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
Supermajorities and faulty agents. Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$, and with this assumption we define $\sigma:=\frac{n+f}{2n}$. Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty agents in $P$. For example, for $f=0$, $\sigma=\frac{1}{2}$; for $f=\frac{1}{2}n$, $\sigma=\frac{3}{4}$; and if everyone but one agent is faulty then $\sigma=\frac{2n-1}{2n}=1-\frac{1}{2n}$, namely unanimity is required. For $n=3f+1$ (the standard assumption), $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
Typically, permissioned consensus protocols wait to hear from “at least nfn-f agents among PP”, which in the standard case of $n=3f+1$ (but not otherwise) happens to be equivalent to “a $\sigma$-supermajority among PP”.
Constitution safety and liveness. A constitution thus defines the protocol’s assumptions about the environment: $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$ and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent; $\Delta\geq\delta$ is an upper bound on the communication time after GST.
Definition 2.2 (Safe and Live Constitution).
A constitution $(P,\sigma,\Delta)$ is safe if $P$ includes at most $f=(2\sigma-1)n$ faulty agents, and live if after GST any message sent among correct agents in $P$ arrives within $\Delta$.
Constitutional Consensus problem definition. A constitutional consensus protocol may have multiple instances. Each instance commences with an initial constitution $(P,\sigma,\Delta)$ that defines its participating agents $P$ and the key protocol parameters $\sigma$ and $\Delta$. As in standard consensus, participants spontaneously issue transactions (these are the inputs to the protocol) and every agent outputs a sequence of transactions.
In addition, constitutional consensus assumes an that participants engage in an unspecified democratic decision process that may result in a decision $d$ to amend the prevailing constitution. The discussion of such democratic processes is deferred to Section 5. For the purpose of the constitutional consensus protocol, all is needed is to assume: (i) the existence of such a process that may result in constitutional amendment decisions, which become known to the participants once made; (ii) that participants can verify a constitutional amendment decisions $d$ to be valid according to the rules specified in Section 5. A valid constitutional amendment decision $d$ defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$, in which the participants $P^{\prime}$ and/or the parameters $\sigma^{\prime}$ and $\Delta^{\prime}$ are the result of amending the prevailing constitution.
In the following, constitution refers to the new constitution of a valid constitutional amendment decision, and an initial constitution is the new constitution of a valid amendment of a constitution with an empty population. Valid amendments are provided to the consensus protocol as inputs by participants that know of them. Thus, the output sequence of a constitutional consensus protocol includes both transactions and valid constitutional amendment decisions, as follows:
Definition 2.3 (Constitutional sequence, epoch, epoch sequence, prevailing constitution, participants).
A constitutional sequence is a sequence $T=t_{1},t_{2},\dots$ of transactions and valid constitutional amendment decisions beginning with an initial constitution. Each constitutional amendment decision $d$ in the sequence defines a new epoch, its epoch sequence is a maximal contiguous subsequence of $T$ that begins with $d$ and does not include another such decision, and its prevailing constitution is the new constitution of $d$. An agent participates in an epoch if it is a member of the population of its prevailing constitution.
The constitutional output sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
Definition 2.4 (Constitutional Consensus Correctness).
In any run of an instance of constitutional consensus in which all valid constitutional amendment decisions made are safe and live, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following safety and liveness properties hold:
  1. 1.
    Safety:
    Consistency
    Every correct agent’s output sequence is a subsequence of $T$.
    Validity
    If a correct agent $p$ outputs a transaction $t$, then $p$ is a participant in $t$’s epoch in $T$.
  2. 2.
    Liveness:
    Epoch Liveness
    If a correct agent $p$ is a participant in an epoch $d$ in $T$, then $p$’s output sequence eventually includes all the transactions in $d$’s epoch sequence. Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in the epoch $d$ is included in $T$.
    Constitution Amendment Liveness
    If a valid constitutional amendment decision $d$ is provided as input to the protocol by a correct agent, then $T$ eventually includes $d$.
Note that if there is no decision to amend the initial constitution then the protocol run has a single epoch with a fixed set of participants. In this case, the definition above coincides with the standard consensus correctness definition.
Note also that progress is guaranteed (after GST) even if the constitution changes infinitely many times. However, normal transactions might starve due to frequent amendments. This definition allows the implementation to prioritize constitution changes over regular transactions, allowing the latter to starve. In practice, amendments are expected to be rare compared to transactions, so we believe this modus operandi to be reasonable.
3 Single Epoch Constitutional Consensus
Here we present the behaviour of the Constitutional Consensus protocol during a single epoch. The protocol is based on the blocklace data structure, which we define in Section 3.1. The protocol’s concepts are detailed in Section 3.2 and its pseudo-code in Section 3.3.
3.1 Blocklace Preliminaries
We use $a\neq b\in X$ as a shorthand for $a\in X\wedge b\in X\wedge a\neq b$.
A $p$-block $b$, with $p\in\Pi$, is a triple $b=(h,x,H)$ where $h$ is the hash of the pair $(x,H)$ signed by $p$, referred to as the identifier of $b$, $x$ is an arbitrary payload, which may be empty, $x=\bot$, in which case we refer to $b$ as an empty block, and $H$ a finite set of block identifiers, the blocks of which are pointed to by $b$. A block with no pointers, $H=\emptyset$, is called initial, in which case it can also be written as a pair $b=(h,x)$. An encoding of a block via a sequence of bits in an agreed-upon form is referred to as a well-formed block. We assume the hash function to be collision-free whp and cryptographic, so that blocks cannot form cycles and we may identify a block $b$ with its identifier $h$. Furthermore, the method of signature allows the public key $p$ to be recovered from a signature by $p$.
A set of blocks $X$ is referred to as a $\sigma$-supermajority given a set of agents $P$ if there is a $\sigma$-supermajority $Q\subseteq P$ such that for every $q\in Q$, $X$ includes at least one $q$-block.
A block $b$ observes itself, any block $b^{\prime}$ pointed to by $b$, and any block observed by $b^{\prime}$. The ‘observes’ relation induces a partial order on any set of blocks. Two blocks that do not observe each other are conflicting, and if by the same agent they are equivocating. A block $b$ approves a block $b^{\prime}$ if $b$ observes $b^{\prime}$ and does not observe any block equivocating with $b^{\prime}$, and a set of blocks $B$ approves $b^{\prime}$ if every $b\in B$ approves $b^{\prime}$.
The closure of a block $b$, denoted $[b]$, is the set of blocks observed by $b$. For a set of blocks $B$, its closure $[B]$ is defined by $B]:=\bigcup_{b\in B}[b]$, which is also the set of blocks observed by $B$. A set of blocks $B$ is closed if $b]\subseteq B$ for every $b\in B$, equivalently if $B=[B]$. A blocklace is a closed set of blocks.
To help avoid confusion, we use $B$ to denote a closed set of blocks (a blocklace) and $D$ or $X$ to denote any set of blocks, not necessarily closed. Also, we refer to a set of blocks by agents in $P\subset\Pi$ as $P$-blocks, the let the $P$-blocks in $B$ denote the set of all blocks in $B$ by agents in $P$.
Given a set of blocks $D$, a pointer to a block $b$ is dangling in $D$ if $b\notin D$. Thus, a blocklace is a set of blocks with no dangling pointers. A block $b\in D$ is a tip of $D$ if no other block $b^{\prime}\neq b\in D$ observes $b$.
The depth of a block $b$, $\textit{depth}(b)$, is 0 if $b$ is initial else $\textit{depth}(b)=\textit{depth}(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$. The depth of a set of blocks $B$ is defined by $\textit{depth}(B):=\max_{b\in B}\textit{depth}(b)$. In the context of blocks, we use most-recent as synonym for deepest. A round $r\geq 0$ in $B$ is the set of all blocks $b\in B$ with $\textit{depth}(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with $\textit{depth}(b)\leq r$, equivalently all rounds in $B$ up to and including $r$. Note that, by definition, blocks of the same round do not observe each other.
3.2 Protocol Concepts
The blocklace created during an epoch begins with a constitutional genesis block $(h,d)$ where $d$ specifies a new constitution $C=(P,\sigma,\Delta)$. In fact $d$ also records the constitutional decision to amend the old constitution (if any) to the new constitution and is signed by multiple agents; these are discussed in Section 5. Here, we focus on the behaviour of the protocol during a single epoch. For now we refer to a blocklace in which all initial blocks are constitutional genesis blocks as a constitutional blocklace. Every participant $p\in P$ commences an epoch $b$ with its blocklace including $b$ and its local epoch identifier $e$ set to $b$, $e\leftarrow b$.
The protocol employs an evolving shared blocklace. Each participant $p\in P$ maintains a local copy of the constitutional blocklace. In the course of an epoch, agents create new blocks that observe the epoch’s genesis block and and disseminate them to other agents. Blocks may include transactions the agents wish to order, if they have any. While a blocklace is a partially ordered set of blocks, the protocol totally orders the blocks in an epoch in a consistent way, producing the same output sequence at all participants in an epoch. Moreover, epochs are also totally ordered, so that the totality of all blocks of all epochs are also totally ordered, even though not always by the same participants
Because correct agents follow the protocol, they do not equivocate. Hence, for a blocklace $B$ created by the protocol with constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
Waves. Given a constitutional blocklace $B$, a wave $W\subseteq B$ in a blocklace is either its constitutional genesis block, referred to as its $0^{th}$-wave, or the set of blocks of depths $r+1$, $r+2$, and $r+3$ in $B$, referred to as the the first, second, and third round of the wave, respectively, for any $r\geq 0$ such that $r\mod 3=0$. Such a wave is the $k^{th}$ wave of $B$, $k:=(r+3)/3$.
For a given $P\subset\Pi$, the function leaderP maps integers to $P$. For each first round of a wave it assigns a formal leader $p\in P$, where a first-round $p$-block $b$ is a leader block if $p=\textit{leader}_{P}(\textit{depth}(b))$. We assume the leader function is defined so that it is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
Finalizing blocks in waves. The consensus protocol commences with a constitutional blocklace and progresses in waves. The protocol strives to finalize a new block in each wave. The first round of a wave includes candidates for finality, the second round includes endorsements of candidates, and the third round includes ratifications.
A first round block is a candidate for finality in two cases—if it is a formal leader block or the only block in its round. These two possibilities typically arise in high- and low-throughput scenarios, respectively. During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with any others, all agents produce second and third round blocks that finalize $b$, in which case $b$ is referred to as spontaneous leader. During high throughput, multiple agents may have transactions to send, and the leader function is used to break symmetry, so that only the formal leader sends a leader block in the first round. Finalized leader blocks of both types induce the total ordering of all blocks, as in other DAG-based protocols [17, 18, 23].
A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows. Note that the next two definitions are mutually-recursive.
Definition 3.1 (Endorse).
Let $b^{\prime}$ be a second round block that approves a first round block $b$ in a wave $W$ in the blocklace $B=[b^{\prime}]$. Let $(P,\sigma)$ be the prevailing constitution.
Then $b^{\prime}$ endorses $b$ if the most-recent wave $W^{\prime}\subset B$ that precedes $b$ is:
  1. 1.
    quiescent and $b$ is the only first-round block approved by $b^{\prime}$, or
  2. 2.
    non-quiescent and $b$ is a $\textit{leader}_{P}$ block approved by $b^{\prime}$.
A set of second-round blocks $X\subset W$ endorses $b$ if every $b^{\prime}\in X$ endorses $b$.
Note that endorsement is an invariant block property, as it depends only on $b^{\prime}]$.
Next, we specify the finality rule. A first-round block $b\in W$ is final if the third round of $W$ has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$. Formally:
Definition 3.2 (Ratified, Final, Ordered, Quiescent).
Assume a blocklace $B$ and a wave $W\subseteq B$. A first-round block $b\in W$ with prevailing constitution $(P,\sigma)$ is:
  1. 1.
    ratified by a third-round block $b^{\prime}\in W$ if $b^{\prime}$ approves a $\sigma$-supermajority in $P$ that endorses $b$, in which case $b$ is also ratified in $W$ and in $B$; and
  2. 2.
    final in $W$ and in $B$ if the third-round of $W$ includes a $\sigma$-supermajority in $P$, referred to as a finalising $\sigma$-supermajority, in which each block ratifies $b$.
The wave $W$ is:
  1. 1.
    finalising if it includes a block $b$ that is final in $W$; and
  2. 2.
    quiescent if, in addition, all blocks in $W$ except perhaps $b$ are empty, and no block in $B$ conflicts with $b$.
In addition, the $0^{th}$-wave is defined to be quiescent and its constitutional genesis block to be final.
Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final. Notice our use of the term “finalised” only applies to such leader blocks; the finalisation of leader blocks induces an order on all the blocks they approve, which are then said to be ordered.
We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$. Note, however, that quiescence is not monotonic, because conflicting and non-empty blocks may be added to a round.
The next observation asserts that a wave’s final block is unique. {observation}[Uniqueness] A wave can have at most one ratified or final block.
Proof 3.3.
The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction.
Quiescence, low- and high-throughput scenarios. The protocol is initially quiescent. It begins operations and continues to operate as long as agents have payloads to send. If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode. It returns to being quiescent following a quiescent wave.
Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block, which is thus finalized by the wave.
Low-throughput good case: The previous wave ending in round $r$ is quiescent, the spontaneous non-empty leader block (red) may be produced at any time after the previous round $r$ has advanced, followed promptly by two rounds of empty blocks, resulting in a finalising and quiescent wave.
High-throughput good case: The formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced and all (or many) of the wave’s blocks are non-empty.
Figure 1: A good-case wave $w$.
We first consider the good-case scenarios summarized in Figure 1. During low-throughput all waves are quiescent. A wave commences when some agent has a new transaction. This agent issues a (first-round) non-empty block, and all agents follow promptly with two rounds of empty blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final, with the new wave also being quiescent.
During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty, with transactions to be ordered by the the next final leader block.
Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own. Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block, with blocks by correct agents to be ordered by the leader of some subsequent finalising wave. Similarly, if during low-throughput there are multiple agents issuing first-round blocks, then typically none of the first-round blocks will become final. In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
To formalize these progress rules, we define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely:
Definition 3.4 (Advanced Round).
Assume a constitutional blocklace $B$ with prevailing constitution $(P,\sigma)$. A round $r$ in a wave $W\subseteq B$ is advanced if it is a:
  1. 1.
    first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a formal leader block or (iii) is preceded by a quiescent wave in $B$ and includes at least one block, or a
  2. 2.
    second or third round that includes blocks by a $\sigma$-supermajority among $P$.
Round 0 is trivially advanced.
Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol. Formally:
Definition 3.5 (Valid Block).
An ordinary block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
Dissemination. The protocol uses two types of dissemination-inducing blocks: nack and inform; other blocks are referred to as ordinary. Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient. Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace.
Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Constitutional Consensus achieves for this model $O(n)$ amortized communication complexity in the good case during high-throughput.
Protocols such as TCP achieve the reliable transfer of a sequence of bytes from the sender to the recipient by employing ack and nack messages for transmission control. The purpose of an ack message is to acknowledge the receipt of a datagram, and of a nack message to report an out-of-order datagram to its sender, e.g., in the case of receiving the $(k+1)^{th}$-datagram before receiving the $k^{th}$-datagram.
In the context of the blocklace, an ‘out-of-order’ block is a block received before some block it points to, and the purpose of a nack-block is indeed to report of such an out-of-order block to its creator (which is not necessarily its sender).
nack blocks are used to achieve reliable broadcast of ordinary blocks. They are needed in order to rectify partial dissemination by Byzantine agents. For example, consider a Byzantine agent that sends a block $b$ to correct agent $p$ but not to another correct agent $q$, and $p$ adds $b$ into its blocklace and sends a block $b^{\prime}$ that depends on $b$. Then when $q$ receives $b^{\prime}$, it cannot add it to its blocklace because its predecessor $b$ is missing. To this end, $q$ sends a nack block to $p$ specifying the blocks it is missing.
Note that out-of-order blocks may naturally arise on a reliable network even during periods of synchrony and even when there are no Byzantine agents, because there is no bound on how quickly messages can arrive — in the example above, $b$’s sender might be correct and $b$ may still be under way to $q$. We therefore introduce a delay before sending a nack. The result is that in the good case, when the network is synchronous and there is no partial-dissemination by Byzantine agents, every correct agent sends its newly created blocks to all other agents and no nack-blocks or other communication is needed.
The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$. If $b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$. These are the tips of $b]\setminus(B\cup\{b\})$. Hence the following definition:
Definition 3.6 (nack-block).
Given a blocklace $B$, a nack-block by $p$ for a block $b=(h,x,H)$ such that $b]\not\subseteq B\cup\{b\}$ is the $p$-block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $b]\setminus(B\cup\{b\})$.
Note that as defined above, $b^{\prime}]\subseteq[b]$, and if $b$ has no pointers dangling in $B$ then a nack-block for $b$ and $B$ is undefined. If the nack-block $b^{\prime}$ is defined, then its pointers $H^{\prime}$ are precisely the non-empty subset of $H$ that is dangling in $B$.
The reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block functions as an ack-block, as it attests to all the blocks known to $p$ at the time of its creation. nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
Definition 3.7 (inform-Block).
Given a blocklace $B$, an inform-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{inform}$ and pointers $H$ to the blocks of depth $r$ in $B$.
If an agent $p$ suspects that the leader of the next round $r+1$ does not proceed since it is missing some round $r$ blocks known to $p$, then $p$ sends the leader an inform-block pointing to the $r$ blocks in its blocklace. If the leader indeed misses any of these blocks, which may happen even after GST due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
3.3 Single-Epoch Constitutional Consensus Pseudocode
In the Constitutional Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$. An epoch identifier $e$ points to the epoch’s constitutional genesis block. We use $B^{e}$ to denote the blocklace of epoch $e$ in $B$, which is the subset of $B$ that observes $e$. The protocol is event-driven. Algorithm 1 describes a single-epoch Constitutional Consensus, in which the prevailing constitution is fixed throughout the run.
Initialisation. The first epoch is initialised with $D\leftarrow\phi$, $B\leftarrow\{b\}$, $e\leftarrow b$, where $b$ is a constitutional genesis block and $\emph{payload}\leftarrow\bot$. The initialisation of subsequent epochs is discussed in Algorithm 2.
Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$. Receive: Upon receipt of a $q$-block $b$, $q\in P$
if $b$ is a nack-block for a $p$-block then judiciously send $[b]$ to $q$,
else if $b$ an inform-block and $b]\not\subseteq B\cup\{b\}$
——- then send once a nack-block for $b$ to $q$,
else if $b$ is a valid ordinary block then $D\leftarrow D\cup\{b\}$.
Accept or Nack: For every $q$-block $b\in D$, $q\in P$,
if $b]\subseteq B\cup\{b\}$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
else if a $b$ was added to $D$ more than $\Delta$ time ago
——- then send once a nack-block for $b^{\prime}$ to $q$.
Issue: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: 1. quiescent and $\emph{payload}\neq\bot$, or 2. non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$. Issue Backlog: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block. Inform Leader: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$. Output: If $B$ contains a final block $b$ of depth greater than the previous final block then output every block in $\tau(b)$ (Definition 3.8) unless it has already been output.
issue a round $k$ block means create a block $b$ with payload and pointers to the $P$-blocks that are tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
Algorithm 1 Constitutional Consensus for a single epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$. Code for agent $p\in P$.
Delays. The protocol makes judicious use of delays, to ensure both liveness and efficiency. Here is the rationale for the various protocol delays, all expressed as multiples of $\Delta$:
$\Delta$ for issuing nack-blocks. In the good case we wish no nack-blocks to be sent at all. Agents wait $\Delta$ before complaining, so if a correct agent sends a block $b$ to $p$ and $q$ at the same time, $q$ may receive $b$ up to $\Delta$ sooner than $q$ and send a block $b^{\prime}$ that depends on $b$ to $p$; but $p$ will receive $b$ within $\Delta$ from receiving $b^{\prime}$. This delay is an optimization to avoid sending nack-blocks in the good case; eliminating it does not hamper safety or liveness.
$2\Delta$ for informing the ensuing round’s leader. In the good case whence all agents are correct after GST, whenever an agent receives a block, all other agents receive it within $\Delta$ time. Thus, once a third round is advanced at the blocklace of a correct non-leader agent $p$, $p$ expects to receive a leader block of the ensuing (first) round within $2\Delta$. If this does not happen, this could be because (i) it is before GST, (ii) the leader is incorrect, or (iii) the leader has not seen yet that the previous round is advanced. This last third case can happen after GST even if the leader is correct, due to Byzantine partial dissemination of third-round blocks. Since $p$ knows that the previous third round is advanced, it sends an inform-block to the leader, which will cause the leader to send a nack-block requesting the blocks it is missing. This $2\Delta$ delay is an optimization to avoid spurious informs in the good case; eliminating it does not hamper safety or liveness.
$9\Delta$ for an unresponsive leader. Following a non-quiescent wave after GST, if the leader $p_{\ell}$ is correct, all correct agents should wait for $p_{\ell}$’s first round block before proceeding to the second round in order to allow that block to be finalized. But because $p_{\ell}$ may be faulty, they need to eventually time-out and stop waiting. To determine the time-out, we consider the worst-case latencies as follows: Assume a third round $r$ is advanced at $p$ at time $t$. If $p$ does not receive a leader block by $t+2\Delta$, it informs $p_{\ell}$, which issues a nack-block, upon receipt of which $p$ sends $p_{\ell}$ any missing blocks. This exchange takes at most $3\Delta$, so by time $t+5\Delta$, $p_{\ell}$ issues a leader block $b$ for round $r+1$, which is received by $p$ by time $t+6\Delta$. Upon receipt of $b$, $p$ might be missing some blocks in $[b]$ and wait $\Delta$ before sending a nack and getting them from $q$ by time $t+9\Delta$. If $p$ does not accept any leader-block by $t+9\Delta$, it concludes that either $p_{\ell}$ is faulty or GST has not arrived yet, and sends a first-round block of its own in order to allow the protocol to proceed to the next round.
Protocol walk-through.
Receive: valid incoming ordinary blocks are inserted to $D$.
Accept or Nack: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
Issue: The variable $r$ always holds the highest advanced round in $B$, and an agent typically issues blocks in round $r+1$. This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced. This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own. The only exception (to sending in round $r+1$) is the Issue Backlog rule: In case $p$ has payload to send, has not sent a round $r$ block, and is not supposed to send a round $r+1$ block, $p$ issues a round $r$ block to ensure its payload is disseminated.
If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block. If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode. During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send. During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader. Here, a long timeout is used ($9\Delta$), before which we try to inform the leader that the round has advanced to give it a chance to progress.
Inform Leader: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
Output: Once the protocol finalizes a block $b$, it orders the blocks approved by $b$ that have not been ordered yet using the function $\tau$. The definition of $\tau$ is essentially-identical to that of Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks. The output of $\tau(b)$ is a sequence that includes all blocks approved by $b$. Formally:
Definition 3.8 ($\tau$).
Assume a safe constitution $(P,\sigma)$ and any fixed topological sort function $\textit{xsort}(b,B,B^{\prime})$ (exclude and sort) that takes a block $b$ and two blocklaces $B\supset B^{\prime}$, and returns a sequence that preserves the ‘observes’ partial order and includes all the non-empty blocks in $B\setminus B^{\prime}$ that are approved by $b$. The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
[…]
When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$. Note that $b^{\prime}$ is not necessarily final in $[b]$, nor final in the closure of the final block with which $\tau$ was called initially. In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
Proposition 3.9.
If the prevailing constitution is safe then $\tau(b)$ does not include equivocations.
Proof 3.10.
Consider an equivocation $b_{1},b_{2}\in[b]$. Consider the first call $\textit{xsort}(b^{\prime},[b^{\prime}],[b^{\prime\prime}])$ for which at least one of $b_{1}$ or $b_{2}$ is in $X:=[b^{\prime}]\setminus[b^{\prime\prime}]$, and the following two cases:
  1. 1.
    Both $b_{1},b_{2}\in X$: Then $b^{\prime}$ approves neither hence neither will be in the output of xsort and hence nor in the output of $\tau$.
  2. 2.
    Only one of them is in $X$, wlog $b_{1}\in X$, and since this is the first such call to xsort it must be that $b_{2}\in[b^{\prime\prime}]$. Then $b^{\prime}$ observes both $b_{1}$ and $b_{2}$ and hence will approve neither, whereas $b^{\prime\prime}$ observes only $b_{2}$, therefore will approve it and therefore only $b_{2}$ will occur in the output of $\tau$, if at all (assuming there are no other equivocations with $b_{2}$).
In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$.
Safety and liveness. We assume that the prevailing constitution is safe and live (Definition 2.2). In Section 4 we prove that $\tau$ is monotonic. This ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor. Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any. In addition, as we shall see, during a protocol run the outputs of calls to $\tau$ by correct agents always occur in the same order, which is the basis for protocol safety.
Liveness is achieved after GST. Whenever there is a post-GST wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block. Therefore, no agent can produce a valid second round block that does not endorse the leader block. So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized. Once any agent has payload to send following a quiescent wave, the agent issues a block, and either the block is finalized in the same wave or the wave is non-quiescent and the protocol switches to high-throughput mode.
Latency and complexity. Regarding latency, we note that after GST and in the good case a leader block (formal or spontaneous) is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
Regarding communication complexity: During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all. In Morpheus [23], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead. Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block. A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block. Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity. To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.
During high-throughput, all blocks carry payloads. If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
4 Safety and Liveness of Single-Epoch Constitutional Consensus
Here we prove the safety and liveness of the single-epoch Constitutional Consensus protocol presented in Algorithm 1. In this case, the members of the founding constitution $C=(P,\sigma,\Delta)$ participate in the protocol for the entire run. The output sequence at every correct participant begins with $C$ and does not include additional constitutions.
4.1 Safety
The safety properties of Definition 2.4 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
{observation}
If the output sequences of every two correct agents are consistent with each other, then single-epoch consensus ensures safety.
Proof 4.1.
Validity trivially holds because every agent participates in every epoch. Assuming the output sequences of correct agents are all consistent with each other, we take $T$ to be their supremum and get Consistency.
Theorem 4.2.
If the founding constitution is safe then single-epoch Constitutional Consensus is safe.
First, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents. We refer to blocklaces and waves satisfying this property as constitutionally-safe wrt $(P,\sigma)$.
For the following proofs we generalize $\tau$ to blocklaces, so that $\tau(B)$ is $\tau(b)$ if $b$ is the most-recent final block in $B$ if there is one, else it is the empty sequence $\Lambda$.
Definition 4.3 ($\tau$ Safety).
A blocklace $B$ is $\tau$-safe if for every wave $W\subset B$ with a block $b$ final in $W$ and any subsequent wave $W^{\prime}\subset B$ with a block $b^{\prime}$ ratified in $W^{\prime}$, $b$ is ratified in $W\cap[b^{\prime}]$.
Proposition 4.4.
A constitutional blocklace with a safe constitution is $\tau$-safe.
Proof 4.5.
Let $B$ be a constitutional blocklace with a safe constitution $(P,\sigma)$, $W\subset B$ a wave with a block $b$ final in $W$, and $W^{\prime}\subset B$ a subsequent wave with a block $b^{\prime}$ ratified in $W^{\prime}$. We have to show that $b$ is ratified in $W\cap[b^{\prime}]$.
Since $b^{\prime}$ is ratified by some block it means that it is valid, since correct agents do not include invalid blocks in their blocklace, let alone endorse them, and a $\sigma$-supermajority includes (a majority of) correct agents. Hence, being a first-round block of a wave, $b^{\prime}$ must observe a $\sigma$-supermajority of the last round of the previous wave, let’s call it the blue supermajority.
Then the blue supermajority and the finalising $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection. Namely, there is a correct agent $p\in P$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $W$ and $W^{\prime}$ are consecutive waves), which in turn must observe a $\sigma$-supermajority of blocks that endorse $b$. Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
Proposition 4.6 ($\tau$ Monotonicity).
$\tau$ is monotonic wrt $\supset$ over constitutionally-safe blocklaces.
Proof 4.7.
We have to show that if $B\subset B^{\prime}$ for blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
By Observation 3.2, all the blocks final in $B$ are also final in $B^{\prime}$. By Observation 3.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth. If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially. Assume $b\neq b^{\prime}$ are the most-recent final blocks in $B,B^{\prime}$, respectively, so necessarily $b^{\prime}$ is more recent than $b$. Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$. By Proposition 4.4 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
{observation}
If the union $B\cup B^{\prime}$ of two blocklaces is constitutionally-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
Proof 4.8.
By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition 4.6), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
Proof 4.9 (Proof of Theorem 4.2).
By assumption, the $P$-blocks in $B\cup B^{\prime}$ include equivocations by at most $f=(2\sigma-1)n$ agents, so this set is constitutionally-safe. Hence, by Observation 4.1, the outputs of any two correct agents running Constitutional Consensus are always consistent. By Observation 4.1, the protocol is safe.
4.2 Liveness
Next, we discuss liveness. Having proved safety, we know that all output sequences of correct agents are subsequences of a common sequence. Let $T$ be the minimal such sequence. Then the set of elements in $T$ is the union of all the elements in output sequences of correct agents. Because all members of $P$ participate in all epochs, Epoch Liveness requires that all their output sequences include all the transactions in $T$, which are all the transactions output by correct agents. Constitution Amendment Liveness trivially holds in the single-epoch case. We conclude with the following observation:
{observation}
[Single-epoch Liveness] A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
Theorem 4.10.
Single-epoch Constitutional Consensus is live.
We first prove that Constitutional Consensus disseminates all correct agents’ blocks and the blocks they depend on.
Proposition 4.11 (Dissemination).
If a block $b$ is issued by a correct agent in Constitutional Consensus then $[b]$ is eventually received by all correct agents.
Proof 4.12.
By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$. Since by assumption $b$ (and possibly other blocks in $b_{p}]$) have not been received by $q$, then after receiving $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$. Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $b_{q}]$, which includes $b$, to $q$. Since the network is reliable, $q$ will eventually received $b$, a contradiction.
The following observation follows directly from the definition of block validity:
{observation}
[Sequential Advance] If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
We next show that the protocol does not get “stuck”, i.e., agents continue to produce blocks when they are needed for progress. Because Constitutional Consensus is quiescent, agents do not have to generate infinitely many blocks. But if some agent issues a block in a round, that round eventually advances.
Proposition 4.13 (Advance).
If a correct agent $p$ sends a block in some round $r$ in Constitutional Consensus, then $r$ eventually becomes advanced at every correct agent.
Proof 4.14.
Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, r1r-1 is advanced in $p_{1}$’s blocklace.
Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$. By Observation 4.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace. By Proposition 4.11, this means that no correct agent sends blocks in rounds >r>r.
Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition 4.11 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
By Proposition 4.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
In case $r$ is a first round following a quiescent wave at $p_{3}$, $r$ advances at $p_{3}$ upon receipt of $b$, and $p_{3}$ issues a round $r+1$ block, a contradiction.
Otherwise, $r$ is the maximum advanced round at $p_{3}$ and is not a first round following a quiescent wave at $p_{3}$. Therefore, $p_{3}$ issues a round $r$ block, a contradiction.
We next prove liveness for finite runs.
Proposition 4.15 (Finite implies quiescence).
Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
Proof 4.16.
Assume by way of contradiction that $p$’s latest advanced round in the run, $r$, is not quiescent. Then, by the Issue rule, $p$ eventually issues a block in round $r+1$ (at most $9\Delta$ time after $r$ is advanced).
Proposition 4.17 (Block issuance).
If an agent $p$ has payload to send, then $p$ eventually issues a block with this payload.
Proof 4.18.
Assume by contradiction that $p$ has a non-empty payload from some time when $p$’s most advanced round is $r_{0}$ onward, and $p$ never issues a block with this payload. First, consider the case that $p$’s rounds do not advance in some suffix of the run. Then by Proposition 4.15, $p$’s blocklace is quiescent. Hence, condition 1 of the Issue rule is satisfied, and the payload is sent. A contradiction. Therefore, $p$’s advanced rounds increase indefinitely. Consider a time when $p$’s highest advanced round $r>r_{0}$ and the payload has not been sent. Then $p$ issues a block either in round $r+1$ (by the Issue rule) or in round $r$ (by the Issue backlog rule). A contradiction.
Proposition 4.19 (Transaction Liveness in finite runs).
Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all input transactions of correct agents are ordered in $B$.
Proof 4.20.
By Proposition 4.11, $B$ includes all ordinary blocks sent by correct agents. By Proposition 4.15, $B$ is quiescent. This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty. Because $b$ is final, $p$ calls $\tau(b)$. Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$. Moreover, by Proposition 4.17, no agents have payloads to send. It follows that all input transactions have been sent in ordinary blocks and ordered.
Proposition 4.21 (Epoch Liveness in finite runs).
Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all transactions output by correct agents are ordered in $B$.
Proof 4.22.
Consider a transaction $t$ output by a correct agent $p_{1}$ and another correct agent $p_{2}$ whose blocklace $B_{2}$ is finite in a suffix of the run. Because $t$ is at some point final in $p_{1}$’s blocklace, its blocklace eventually includes a block $b$ with $t$ in its payload and a super-majority of blocks that observe $b$. This super-majority includes at least one $p_{3}$-block $b^{\prime}$ by a correct agent $p_{3}$. By Proposition 4.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$. Because $B_{2}$ is finite in a suffix of the run, by Proposition 4.15, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed.
Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalising. Because quiescence is not monotonic, we consider here only waves $k$ for which the predecessor wave, k1k-1, is not quiescent at any correct agent at any time.
Proposition 4.23 (First round progress after GST).
Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt>GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time. Then $p$ sends a block in round $r$ if and only if it is the round’s leader. Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
Proof 4.24.
Let $p_{\ell}$ be the correct leader of round $r$. If $p=p_{\ell}$ then once r1r-1 is advanced, $p$ immediately issues a leader block for round $r$, causing round $r$ to advance, and issues a block endorsing it in round $r+1$.
Otherwise, $p$ waits for a block from $p_{\ell}$ until time $t+2\Delta$ and then sends $p_{\ell}$ an inform-block, which is received by $t+3\Delta$, causing $p_{\ell}$ to send $p$ a nack-block and get the blocks it is missing by time $t+5\Delta$. Because round r1r-1 is advanced and its wave is not quiescent, $p_{\ell}$ sends a leader block, which is received by $p$ by time $t+6\Delta$. Some predecessors of this block might be missing, causing $p$ to wait $\Delta$ and then send a nack-block and get the missing blocks allowing it to accept the leader block by time $t+9\Delta$.
When the leader block is received, the round advances and $p$’s round $r+1$ block endorses the leader block.
Below, we say that a round $r$ advances at time $t$ if there is some correct agent $p$ such that round $r$ is advanced in $p$’s blocklace at time $t$ and $r$ is not advanced at any correct agent before time $t$.
Proposition 4.25 (Progress after GST).
Let $r$ be a first round with a correct leader s.t. round r1r-1 advances after GST and the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time. Then a round $r$ leader block is finalized in its wave at all correct agents.
Proof 4.26.
By Proposition 4.23, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$. This means that it is impossible to create a valid round $r+1$ block that does not endorse the leader, as such a block would have to point to a super-majority of round $r$ blocks excluding the leader block. Therefore, every valid round $r+2$ block ratifies the leader block, and the block is finalized in round $r+2$ as soon as it advances at any correct agent.
Next, we consider runs with infinitely many quiescent waves.
Proposition 4.27 (Infinite quiescent waves).
Consider a run of Constitutional Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. Then all the blocks of correct agents are ordered by all agents in this run.
Proof 4.28.
Let $B$ be the (infinite) union of all blocklaces in the run. Let $b_{1}$ be a block of a correct agent in $B$. By Proposition 4.11, $b_{1}$ is eventually included in all correct agents’ blocklaces. Therefore, it is approved by all correct agents’ blocks from some wave onward. Because every quiescent wave is finalising, and finalization is monotonic, there are infinitely many finalising waves in $B$. Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$. Once $b_{2}$ is final at some agent’s blocklace, $b_{1}$ is ordered at that agent. We will show that $b_{2}$ is finalized at all agents.
Observe that since $b_{2}$ is finalized in wave $k$, it is approved by a super-majority in wave $k$. Because every valid first round block approves a super-majority in the previous round and there is at least one correct agent at the intersection of any two super-majorities, every valid first round block in any round $k^{\prime}>k$ observes $b_{2}$.
Consider a finalising wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$. Then $b_{3}$ is endorsed by a super-majority in the second round. Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 4.11). If $b_{3}$ is not a leader block, then to be validly endorsed, $b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
By the same token, if any non-leader first round block $b$ is endorsed by a valid block of any agent in the second round of wave $k+1$, then $[b]$ is quiescent. This, in turn, means that every third round block in the wave either ratifies the leader block or observes a block whose closure is quiescent.
So when the third round of wave $k+1$ advances, every correct agent either observes a super-majority of blocks that ratify the leader block or observes a block whose closure is quiescent. In the former case the leader block is finalized, ordering all preceding ordinary blocks including $b_{2}$, and in the latter $b_{2}$ is already final in the block’s closure.
We are now ready to prove the liveness theorem.
Proof 4.29 (Proof of Theorem 4.10).
By Observation 4.2, it suffices to show that every input or output transaction of a correct agent is eventually included in the output sequence of every correct agent. By Proposition 4.17, all input transactions in agents’ payloads are eventually sent in (ordinary) blocks. Similarly, every output transaction of a correct agent is included in some ordinary block. Let $b$ be a non-empty ordinary $p$-block included in some correct agent’s blocklace in a run of Constitutional Consensus. If $p$ is correct, then by Proposition 4.11, every correct agent receives and incorporates $b$ in its local blocklace. Otherwise, $b$ is finalized in some blocklace, implying that it is approved by at least one $q$-block $b^{\prime}$ by a correct agent $q$. Again, by Proposition 4.11, every correct agent receives and incorporates every block in $b^{\prime}]$, including $b$, in its local blocklace.
We consider three cases: First, assume the run is finite. By Propositions 4.19 and 4.21, all non-empty ordinary blocks sent or output by correct agents in this run are ordered at all correct agents, and the theorem holds. Second, consider a run in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. By Proposition 4.27, if $p$ is correct then $b$ is ordered, and otherwise $b^{\prime}$ is ordered, implying that $b$ is ordered as well because $b\in[b^{\prime}]$.
Finally, assume the run is infinite and contains only finitely many waves that are quiescent in at least one correct agent’s blocklace at some point in time. Let $k$ be a wave whose last round advances after GST and after which no wave is quiescent at any correct agent’s blocklace. By fairness of leader selection, there are inifinitely many waves $k^{\prime}>k$ with a correct leader. By Proposition 4.25, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds.
5 Constitutional Consensus with Constitutional Amendments
The problem of reconfiguration of a consensus protocol—namely changing the set of its participants—has engaged the distributed computing research community for decades [40, 42, 2, 40], ever since it was first formulated in Lamport’s Paxos paper [21]. Here we explore the more general problem of constitutional amendments, explained and defined below. Supporting constitutional amendments makes the consensus protocol grassroots, as proven in Section 6.
5.1 Constitutional Consensus: From General to Specific
In its most general sense, a constitutional consensus protocol is provided with an initial constitution that specifies an initial population, an initial consensus protocol, a rule for amending both, and a rule for amending the constitution itself. The notion of constitutional amendment is different from the well-studied notion of reconfiguration in at least two respects:
  1. 1.
    Reconfiguration typically refers to changing the set of active agents (servers, processes), whereas a constitution of a consensus protocol may address any aspect of it, including the agents executing it, but also the protocol itself and its parameters.
  2. 2.
    Reconfiguration studies are not concerned with who makes reconfiguration decisions: They are typically assumed to descend on obedient agents “out of the blue”. Amending the constitution of a protocol, on the other hand, is done by the volitional agents (e.g., people) executing the protocol, in accordance with the prevailing constitutional amendment rule.
In this work we consider a restricted type of constitution $(P,\sigma,\Delta)$ (Definition 2.1), the components of which affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony. Even more so, they are key to Constitutional Consensus, given that its primary aim is to execute digital social contracts [9, 38]:
  1. 1.
    $P$: The people $P$ are the parties to the contract—members of the democratic digital community, community bank, cooperative, etc. Clearly, they should be sovereign to accept new members and remove existing members, namely to amend $P$.
  2. 2.
    $\sigma$: As the level of knowledge and trust among the people $P$ evolves, possibly as a result of amending $P$, so should the $\sigma$-supermajority among $P$ be amended to reflect that: If the people in $P$ conclude that they can trust each other more than initially anticipated, $\sigma$ could be decreased. Adding new people to $P$, initially known to and trusted by only a fraction of $P$, may entail increasing $\sigma$, even if temporarily, in order to ensure that sufficiently many trusted agents are part of every decision.
  3. 3.
    $\Delta$: The delay parameter $\Delta$ is also highly-sensitive to the composition of the population $P$: If all have high-performance smartphones that are well-connected, $\Delta$ can be decreased; if new members with lower-capacity devices are admitted, $\Delta$ may need to be increased. In addition, technology advances may allow decreasing $\Delta$. Thus, the prevailing population $P$ should be sovereign to amend the value $\Delta$ as they see fit.
5.2 Democratic Constitutional Amendments
The Constitutional Consensus protocol attempts to implement valid constitutional amendment decisions as soon as they are made, but is independent of how these decisions are made. Computational Social Choice research [6] has investigated democratic decision-making processes extensively. Here we describe possible democratic processes for amending each of the components of a constitution $(P,\sigma,\Delta)$.
(1) Amending $P$. For amending $P$ to $P^{\prime}$, we require that a $\sigma$-supermajority among $P$ and that all new members of $P^{\prime}$, namely those in $P^{\prime}\setminus P$, approve the decision. The purpose of the supermajority is to ensure that a majority of the correct agents in $P$ are willing to hand-over control to $P^{\prime}$. The purpose of requiring all new participants in $P^{\prime}$ to sign is, first, to respect their autonomy—an agent should agree to participate—and, second, to avoid the ‘dead sailors problem’ [21], in which all of $P^{\prime}$ are new but in fact dead, resulting in a constitutional stalemate.
The respect for autonomy of new participants has a practical corollary: An honest person (who would participate as a correct agent) would not join a protocol instance if they are concerned that the prevailing population may have too many Byzantines, namely that the prevailing constitution is not safe. The flip side of this concern—that the prevailing population not add malicious agents—has been studied extensively, as discussed next.
A constitutional consensus protocol shares a fundamental challenge with permissionless consensus protocols: Both allow the participating agent set to change dynamically, yet must prevent Sybil attacks where a single entity creates multiple digital agents to gain disproportionate influence. Permissionless protocols address this by requiring agents to prove control of scarce resources—computational power in Proof-of-Work [27] or financial stake in Proof-of-Stake [19, 8]. We expect realization of constitutional consensus to take a people’s-based approach: To rely on trust relations among real people, and admit new agents only when existing members know, trust, and verify them as genuine individuals. This requires (i) sybil-resistant admission protocols [28, 33] during community formation to ensure that digital agents genuinely represent distinct people rather than fake or duplicate identities and (ii) sybil-resilient democratic governance [32, 24], to ensure that sybils cannot overpower the genuine agents in democratic decisions. These mechanisms are complementary to this work.
(2) Amending $\sigma$. Amending the supermajority by which a constitution can be amended has been studied by Abramowitz et al. [1], who showed that the $h$-rule (similar to the $h$-index of author citation metrics) is the only rule that satisfies certain intuitive and self-evident axioms. According to the $h$-rule, each agent in the prevailing population $P$ votes by stating their most-preferred new value of $\sigma$, in a process left unspecified. The value of $\sigma$ is increased to the maximal $\sigma^{\prime}>\sigma$ for which there is a $\sigma^{\prime}$-supermajority among the prevailing population $P$ who voted for a value greater or equal to $\sigma^{\prime}$. Else $\sigma$ is decreased to the minimal $\sigma^{\prime}$ for which there is a $\sigma$-supermajority among the prevailing population $P$ who voted for a value less than or equal to $\sigma^{\prime}$. Else $\sigma$ remains unchanged.
Thus, to change $\sigma$ from $\frac{2}{3}$ to $\frac{3}{4}$, the approval of a $\frac{3}{4}$-supermajority among $P$ is needed, and to change to unanimity ($\sigma=\frac{2|P|-1}{2|P|}$), approval by all members of $P$ is needed. And if the prevailing constitution states unanimity, then changing back to a $\frac{2}{3}$-supermajority requires unanimous consent among $P$.
The $h$-rule ensures that changing the supermajority threshold requires meeting that very threshold: To raise $\sigma$ to $\sigma^{\prime}$, you need a $\sigma^{\prime}$-supermajority supporting it (preventing minorities from imposing stricter requirements on the majority), while to lower $\sigma$ from the current value, you need the current $\sigma$-supermajority to agree (preventing minorities from weakening the constitution’s protection).
(3) Amending $\Delta$. The value of the parameter $\Delta$ should be increased or decreased only if the majority of the correct agents wish so. To this end, we offer the Suppress Outer-f parameter update rule [^1] In reference [32] the rule refers to ‘outer-$\sigma$’, however the definition of $\sigma$ there is $\frac{f}{n}$, so we changed the name here to avoid confusion. by Shahaf et al. [32], as follows: Each member of the prevailing population $P$ votes by stating their most-preferred new value of $\Delta$, through a process left unspecified. The vote on $\Delta$ should not be based on the subjective experience of each member of $P$, but rather on the perception of each member of $P$ of the “common good” [29].
If the median of the votes is larger than $\Delta$, then the maximal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}>\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$. Symmetrically, if the median of the votes is smaller than $\Delta$, then the minimal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}<\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$. Else $\Delta$ remains unchanged.
We note several desirable properties of the Suppress Outer-f rule: If all Byzantines wish to increase $\Delta$ to slow-down the protocol, and the correct agents do not, then $\Delta$ will not be increased. Similarly, if all Byzantines wish to decrease $\Delta$ to undermine the liveness of the protocol, and the correct agents do not, then $\Delta$ will not be decreased. If all correct agents unanimously wish to update $\Delta$ to $\Delta^{\prime}$ then $\Delta$ will change to $\Delta^{\prime}$ no matter what the Byzantine agents vote, provided $f<\frac{1}{3}n$. In summary:
A constitution $(P,\sigma,\Delta)$ could be amended democratically to $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$ as follows: 1. $P$ to $P^{\prime}$ via the approval of both a $\sigma$-supermajority among $P$ and all members of $P^{\prime}\setminus P$. 2. $\sigma$ to $\sigma^{\prime}$ via the h-rule for constitutional amendment [1]. 3. $\Delta$ to $\Delta^{\prime}$ via the suppress outer-f parameter update rule [32].
5.3 Constitutional Amendments
While we suggest above democratic decision-making processes for amending the three components of the constitution, their specifics are immaterial to the Constitutional Consensus protocol. The only thing that matters is the interface between the decision making process and the protocol, which takes the form of constitutional amendment decisions. The following definition assumes a Constitutional Consensus protocol instance identifier Id; we discuss its selection below.
Definition 5.1 (Constitutional Amendment Decision).
A constitutional amendment decision is a pair $d=(h,x)$, where $x=(\textsc{amend},\textit{Id},i,C,C^{\prime})$, Id is an instance identifier, $i\geq 1$ is the index of the decision within this instance, $C,C^{\prime}$ are constitutions, referred to as the old and new constitutions of $d$, respectively, and $h$ is a multisignature of the hash of $x$ by some $Q\subset\Pi$. If $i=1$ then $C=\bot$ and $P_{1}$ are referred to as the founders of the protocol instance Id—they are the signatories of its initial constitution and the participants of the first epoch.
Note that the structure of $d$ is that of a constitutional genesis block. During a run of a protocol instance Id, agents produce a possibly-infinite ordered sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$ using democratic decision-making processes discussed above or others, where $d_{i}=(h_{i},x_{i})$ and $x_{i}=(\textsc{amend},\textit{Id},i,C_{i-1},C_{i})$. The protocol commences with $d_{1}$ as the constitutional genesis block of the first epoch, and with each subsequent $d_{i}$ serving as the constitutional genesis block of the $i^{th}$ epoch.
Definition 5.2 (Valid Constitutional Amendment Decision and Block).
Given a sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$, the amendment $d_{i}$ with prevailing constitution $C_{i-1}=(P_{i-1},\sigma_{i-1},\Delta_{i-1})$, new constitution $C_{i}=(P_{i},\sigma_{i},\Delta_{i})$, and signatories $Q_{i}$ is valid if $i=1$, $P_{0}=\emptyset$ and $Q_{1}=P_{1}$, or $i>1$, $d_{i-1}$ is valid and the signatories $Q_{i}$ are (i) a $\sigma_{i-1}$-supermajority in $P_{i-1}$, (ii) a $\sigma_{i}$-supermajority in $P_{i}$, and (iii) $(P_{i}\setminus P_{i-1})\subseteq Q_{i}$. The sequence is valid if all its members are valid. Two non-identical constitutional amendment decisions are conflicting if they have the same instance identifiers and the same index. Given a sequence of valid constitutional amendment decisions, a block is constitutional if it includes a valid constitutional amendment decision as payload.
Note that for agents to determine that a constitutional amendment decision is valid, they need access to all its predecessors. Also note that any two consecutive supermajorities may overlap, even be identical. In particular, if only $\Delta_{i}$ changes, a $\sigma_{i}$-supermajority among $P_{i}$ is needed. If only $\sigma_{i}$ changes then a single $\max(\sigma_{i-1},\sigma_{i})$-supermajority among $P_{i}$ is needed; this ensures that the $h$-rule [1] is followed. If only $P_{i}$ changes then a $\sigma_{i}$-supermajority among $P_{i-1}$, a $\sigma_{i}$-supermajority among $P_{i}$, and all new members in $P_{i-1}$, are needed.
We assume that correct agents do not sign invalid or two conflicting constitutional amendments, and hence can make the following observation.
{observation}
[Safety of Constitutional Amendments] Given a valid finite sequence of constitutional amendment decisions $D=d_{1},\ldots d_{i}$, $i\geq 1$, if every $C_{i}$ in $D$ is safe, then the constitutional amendment process cannot extend $D$ with two conflicting valid constitutional amendments.
Proof 5.3.
By a counting argument, for this to happen a correct agent would have to sign two conflicting constitutional amendments, a contradiction.
Note that while the safety depends only on $\sigma_{i}$ and $P_{i}$, validity depends also on $\sigma_{i-1}$ and $P_{i-1}$, in order to adhere to the $h$-rule.
Grassroots selection of protocol instance identifier. Each constitutional amendment decision is associated with an instance Id of the Constitutional Consensus protocol. In a grassroots setup, the identifier of each instance should be unique in order to allow for multiple instances to operate concurrently without interference, even if they share some or all of the agents, and should be chosen without central authority, central resource, or coordination among disjoint instances. This can be achieved, for example, by having the identifier Id of a protocol instance consist of (the hash of) the founders (initial population) of the instance, each paired with the maximal serial number of any Constitutional Consensus protocol instance they have founded. This ensures Id to be unique if the founders include at least one correct agent, who would not sign the initial constitutional decision if its Id is not formed correctly.
Epochs and agent responsiveness. The standard assumption in consensus protocols is that all correct agents are responsive throughout the run. However, this assumption is unreasonable in a protocol like Constitutional Consensus in which the set of participating agents changes dynamically, especially since the set $\Pi$ of potential participants is infinite. Hence, we only assume that a correct agent $p$ is responsive only during an epoch in which it participates. The one exception is the case of agents that are behind, still in “older” epochs. To complete the epoch they may need to receive such blocks from former correct participants in that epoch. Hence, agent may still need to respond to nack-blocks pertaining to blocks they have issued in past epoch.
Still, an agent may determine that en epoch is ‘closed’ (and no further nack-blocks need to be served from it) if it has evidence that all participants in the epoch have ended it. In particular, an epoch can be ‘closed’ when all its participants have coronated the constitutional amendment decision that ended (see End Epoch below).
Note that the interaction in which a non-responsive agent $p$ needs to be ‘woken up’ in order to sign a constitutional amendment in which $p$ is a new participant happens outside the Constitutional Consensus protocol. In the practical settings we envision this would happen through social interactions (in-person or digital) between existing participants and prospective new participants.
5.4 Epoch Transitions via Constitutional Amendments
Protocol concepts. Every epoch commences with a constitutional genesis block that specifies its prevailing constitution. If, during the epoch, members of the prevailing population become aware of a valid decision $d$ to amend the prevailing constitution to a new constitution, they cease normal operation and only endeavour to finalize this decision. Once an agent has a wave $W$ that finalizes a constitutional block with $d$, it coronates $d$ by producing a coronate block with payload $(\textsc{coronate},d)$, sending it to both the old and new populations of $d$, and ceasing participation in the epoch. Members of the old population that receive it and have yet to coronate $d$ can nack any missing blocks to obtain $W$ and coronate $d$ themselves. A member of the new population that receives a $\sigma$-supermajority of coronate blocks may commence the new epoch with $d$ as the constitutional genesis block. An agent that is a member of both the new and old populations commences the new epoch after both receiving a $\sigma$-supermajority of coronate blocks and sending one of its own (implying that $d$ is final in its blocklace).
The protocol. The protocol produces blocks also with constitutional amendment decisions as payload; such a block is referred to as a constitutional block. When $\tau$ orders a constitutional block with amendment $d$ for the first time, the protocol outputs the new constitution. Subsequent ordered blocks with the same amendment are suppressed from the output.
The protocol is extended in Algorithm 2 as follows:
  1. 1.
    Amendment Finalization occurs whenever the agent $p$ knows of a valid constitutional amendment $d$ that has not been finalized, upon which it continues following the protocol but only producing constitutional blocks with $d$ as a payload, until $d$ is finalized. Here, knows has a dual meaning: Either $p$ knows that the external democratic decision making process resulted in a new valid constitutional amendment $d$, or $p$ has a constitutional block with $d$. In either case, no constitutional block with $d$ has been finalized yet in $B$.
  2. 2.
    End Epoch applies when a constitutional block with $d$ is finalised, upon which members of the old population create a coronate block with $d$, send it to all members of the old and new populations, and cease participation in the old epoch. Coronation also plays a role in the democratic decision process: A correct agent would not sign a new constitutional amendment decision before coronating the last constitutional amendment it was a signatory to.
  3. 3.
    Start Epoch applies when any agent $p$ in the new population receives from the old population a supermajority of coronate block with decision $d$ (and provided $p$ issued a coronate block if it is also a member of the old population), upon which $p$ creates a constitutional genesis block $b$ with $d$ and updates its epoch identifier to be $b$.
  4. 4.
    Respond to Belated nack applies when any agent $p$ receives a nack $q$-block $b$ for a $p$-block of a past epoch $e$, upon which $p$ responds to the nack to help $q$ complete this past epoch. Note that if $p$ had previously received a coronate $q$-block ending $e$, then $p$ knows that $q$ is not missing any epoch $e$ blocks; in this case “judicially send…” would be a no-op. Moreover, an agent may ‘write off’ an epoch $e$ as complete after receiving from all participants in $e$ blocks coronating the end of $e$.
Amendment Finalization by $p$: Upon issuing an ordinary block, if $p$ knows of a valid constitutional amendment decision $d$ that is not final in $B$ then issue the block with $d$ instead of with payload. End Epoch by $p$: Upon finalising in $B$ a constitutional block $b$ with constitutional amendment $d$, prevailing population $P$, and new population $P^{\prime}$, call $\tau(b)$, create a $(\textsc{coronate},d)$ block, send it to members of $P\cup P^{\prime}$, and cease participation in the epoch. Start Epoch by $q$: Let $d$ be a valid decision to amend the prevailing constitution, with new population $P^{\prime}$.
Upon receipt by $q\in P^{\prime}$ of a $\sigma$-supermajority among $P$ of $(\textsc{coronate},d)$ blocks, and provided $q$ previously issued such a block if $q\in P$, commence participation in the new epoch by $e\leftarrow d$ as the new epoch constitutional genesis block, $B\leftarrow B\cup\{d\}$, and retaining $D$ and payload.
Respond to Belated nack by $q$: Upon receipt of a nack $q^{\prime}$-block for a $q$-block $b$ of a past epoch $e$ for which $q^{\prime}\in\Pi$ is a participant, judiciously send $[b]$ to $q^{\prime}$.
Algorithm 2 Incorporating Constitutional Amendments in Constitutional Consensus.
Code for any participant $p$ in the prevailing constitution and for any $q\in\Pi$.
In the good case, a new constitutional amendment $d$ is issued by the leader of a wave and finalized within that same wave. In the worst case, it might take arbitrarily many waves to finalize $d$ prior to GST or due to faulty leaders. After GST, a constitutional amendment block issued by a correct leader among $P$ is bound to be finalized (and such a leader is bound to be selected eventually due to the fairness of $\textit{leader}_{P}$). Once finalized, correct members of $P$ will issue a coronate block and correct members of $P^{\prime}$ will receive it, allowing the current epoch to end and the next one to begin. This argument informally establishes the liveness of the transition among constitutions due to constitutional amendments.
5.5 Safety and Liveness of Constitutional Amendments
Theorem 5.4 (Constitutional Amendment Correctness).
The Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) is correct.
We need to prove that in every run of the protocol, there exists a sequence $T$ s.t. the two safety properties and three liveness properties of Definition 2.4 hold w.r.t. $T$. Note that the $i^{th}$ epoch $e_{i}$ is in fact $d_{i}$. We have proven that Algorithm 1 is correct in each epoch. Therefore, there is a sequence $T^{i}$ so that the output sequences produced by all participants in epoch $e_{i}$ are consistent and valid w.r.t. $T^{i}$.
Let $T$ be the concatenation of all these sequences, $T:=T^{1}\cdot T^{2}\cdots$. We next prove that the protocol satisfies safety and liveness properties w.r.t. $T$. The theorem will follow.
Lemma 5.5 (Constitutional Amendment Safety).
Every run of the Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) satisfies safety w.r.t. $T$ defined above.
Proof 5.6.
Validity immediately follows from the fact that agents participate only in epochs of which they are members. Consistency follows from the fact that agents participate in epochs sequentially, namely, they stop running the protocol in epoch $i$ before starting epoch $j$.
Lemma 5.7 (Constitutional Amendment Liveness).
Every run of the Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) satisfies liveness w.r.t. $T$ defined above.
Proof 5.8.
Consider an epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$, and the two liveness requirements. If $e$ is the last epoch then Epoch Liveness follows from the liveness of the single-epoch Algorithm 1, Theorem 4.10. Else assume $e$ is not the last epoch, and therefore it must have ended with a coronation by a $\sigma$-supermajority among $P$, all observing a wave finalizing a constitutional block $b$ with payload $d$.
Let $p\in P$ be a correct agent. We argue that $p$ must eventually coronate $d$, and while doing so, call $\tau(b)$. As all correct participants in the epoch $e$ end it by calling $\tau(b)$, all would output the same blocks for $e$. If $p$ is among the $\sigma$-supermajority coronating $d$ we are done. Else there is at least one correct agent $q\in P$ that coronated $d$ and sent to $p$ a coronate block $b^{\prime}$ observing a wave $W$ that finalizes $b$ . Since $q$ and $p$ are correct, eventually a $\textsc{nack}(b^{\prime})$-block by $q$ will be received by $p$, who will respond (even after the epoch $e$ has ended, thanks to the Respond to Belated nack rule) by sending to $q$ any blocks in $W$ and their predecessors that $q$ is missing, upon receipt of which $p$ would issue a $(\textsc{coronate},d)$ block and output $\tau(b)$.
Considering Constitution Amendment Liveness, once a constitutional amendment $d$ is known, all correct participants follow the high-throughput protocol and produce only blocks with payload $d$. Eventually a correct leader will finalize its block, following the same argument as in the proof of Theorem 4.10, upon which all correct participants will eventually issue $(\textsc{coronate},d)$ blocks.
6 Constitutional Consensus is Grassroots
We recall the formal definition of a grassroots protocol from reference [38], instantiate it informally (bypassing the formalities of the definition in [38]) to the Constitutional Consensus protocol, and then argue that the protocol is grassroots.
Definition 6.1 (Oblivious, Interactive, Grassroots).
A protocol $\mathcal{F}$ is:
  1. 1.
    oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, a run of the protocol over $P$ can proceed just the same in the presence of agents in $P^{\prime}$.
  2. 2.
    interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every run of the protocol over $P$, the protocol can proceed from any configuration $c$ in $r$ to interact with members in $P^{\prime}\setminus P$, namely produce a run that is not possible only over $P$.
  3. 3.
    grassroots if it is oblivious and interactive.
Note that the definition is stated in terms of any two populations $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$.
Being oblivious, in the context of Constitutional Consensus, means that if the prevailing populations of the protocol during a run are always included in $P$, then the presence of additional agents hanging around, those in $P^{\prime}\setminus P$, which do not participate in the protocol, should not interfere with the protocol or affect its behaviour. Clearly, this is the case with Constitutional Consensus, as its prevailing populations may always choose not to amend the constitution to include members outside $P$. Hence Constitutional Consensus satisfies the requirement of being oblivious.
Being interactive, in the context of Constitutional Consensus, means that if the prevailing populations of the protocol in some prefix of a run ending in a configuration $c$ are included in a set $P$, then the protocol has a computation starting from $c$ that interacts with members outside $P$ in a meaningful way, namely the protocol has a computation from $c$ that reaches a configuration $c^{\prime}$, in which members of $P^{\prime}\setminus P$ are included in the prevailing population, and thus $c^{\prime}$ cannot be reached if the prevailing populations are all restricted to $P$. Clearly, members of the prevailing population $P$, starting from configuration $c$, may amend the constitution to include in $P$ a new member $p^{\prime}\in(P^{\prime}\setminus P)$ and thus reach a configuration $c^{\prime}$ in which $p^{\prime}\notin P$ is a member. Hence Constitutional Consensus satisfies the requirement of being interactive. Together, these two properties imply that Constitutional Consensus is indeed grassroots. We thus conclude:
Theorem 6.2.
Constitutional Consensus is grassroots.
Note that the two requirements cannot even be expressed without in the context of the basic (no constitutional amendments) Constitutional Consensus protocol, as the population $P$ is given and fixed. Also note that Constitutional Consensus being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
7 Conclusions
The Constitutional Consensus protocol and its extensions are novel in several respects, all devised in order to serve the needs of future grassroots platforms that operate digital social contracts. Initially, we expect low-throughput applications to prevail, as small digital communities engaged in constitutional democratic conduct use the platform to grow and evolve [39]. Even small community banks and cooperatives [37] may still be, at least initially, on the low-throughput side. As communities federate to form ever-larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow. Until then, simulating the various scenarios would be productive future work.